# Leveraging Graph Neural Networks and Multi-Agent Reinforcement Learning for Inventory Control in Supply Chains

### v1

@@ Abstract
Our approach redefines the action space by parameterizing heuristic inventory control policies, making it adaptive as the parameters dynamically adjust based on system conditions. By leveraging the inherent graph structure of supply chains, our framework enables agents to learn the system’s topology, and we employ a centralized learning, decentralized execution scheme that allows agents to learn collaboratively while overcoming information-sharing constraints. Additionally, we incorporate global mean pooling and regularization techniques to enhance performance.
@@ 1.1 Inventory Control
The theory of inventory control can be traced back to the news-vendor problem (Edgeworth, 1888; Clark and Scarf, 1960), and the first widely used numerical solutions for inventory optimization seems to be the Economic Order Quantity (EOQ) model (Erlenkotter, 1990). These works were fundamental to the widely known policies in inventory control theory: (R,S), (s,S), (R,s,S) and (R,Q) (Federgruen and Zheng, 1992; Silver et al., 1998). Over the years, control theory has had an influence on production planning and inventory control (Aggarwal, 1974; Jackson et al., 2020). The complexity of modern systems has driven dynamic programming applications, unfortunately, obtaining exact analytical solutions often proves infeasible due to computational demands (Bellman, 1952). This has fuelled the adoption of advanced techniques beyond traditional closed-form solutions. On one hand, stochastic (Küçükyavuz, 2011) and distributionally-robust optimization approaches (Bertsimas et al., 2019; Qiu et al., 2021), while promising, face challenges such as online tractability, limiting their practical applicability in real-world supply chain scenarios. On the other hand, traditional approaches leveraging heuristics, while offering a simpler implementation, struggle to adapt to the dynamic and complex nature of modern supply chains.
@@ 1.2 Reinforcement learning for Inventory Control
Reinforcement learning (RL) has emerged as a promising alternative for addressing the challenges of stochastic sequential decision-making problems. Its ability to excel in complex, dynamic environments and handle uncertainty, places RL as a valuable tool to enhance decision making in supply chains. While closely linked to dynamic programming, RL offers a general solution to identify approximate optimal policies for stochastic processes by leveraging the Bellman optimality equation to iteratively update value functions and improve decision-making policies over time. RL also provides a cost-effective solution for decision-making by enabling offline training, which reduces the online computational overhead compared to optimization approaches that require continual updates in receding or shrinking horizon frameworks. However, RL’s effectiveness relies on the availability of information sharing among the entities within the supply chain. In scenarios where multiple interconnected entities are involved, ensuring seamless information exchange can be challenging due to factors such as data privacy concerns, proprietary information, or communication constraints. This limitation may hinder the adoption and implementation of RL-based approaches in complex supply chain environments, highlighting the need for strategies to address information sharing challenges effectively.
@@ 1.3 Multi-Agent Reinforcement Learning
This challenge in information sharing motivates the exploration of multi-agent reinforcement learning (MARL) approaches in supply chain management. MARL frameworks allow multiple entities within a supply chain to autonomously learn and adapt their decision-making processes while interacting with each other. By leveraging MARL, supply chain entities can collaborate and coordinate their actions more effectively, leading to improved overall performance and resilience against disruptions. Additionally, MARL provides a scalable approach to handle the complexities of large-scale supply chain systems by distributing decision-making across multiple agents.
However, despite its potential, MARL faces several challenges. One issue is the non-stationarity of the learning environment inherent in MARL (Tan, 1993; Tampuu et al., 2017). This occurs when multiple agents are learning simultaneously so the transition dynamics are not stationary. One setting commonly used to overcome this is to use a central critic during training that has access to global observations and actions (Nekoei et al., 2023). This is known as the Centralised Training, Decentralized Execution learning paradigm. However, while adding more information can help mitigate the non-stationarity problem, it can also lead to new issues. Naively concatenating all available information can result in information overload and inefficiencies (Lowe et al., 2017; Yu et al., 2022; Nayak et al., 2023). This indiscriminate accumulation of data can cause policy overfitting, where agents develop strategies that perform well on the excessive training information but poorly in real-world scenarios due to the lack of generalization (Nayak et al., 2023; Hu et al., 2021). Therefore, the development of novel techniques for smart information aggregation is required to avoid policy overfitting and enhance the efficacy of MARL frameworks in supply chain management.
1.4 Related Work
Given the inherent complexity and uncertainty in inventory management systems, there are several approaches that address specific challenges within this domain. In this section, we highlight some related previous work and methodologies tackling the inventory management problem.
One example to find optimal inventory control policies is using traditional heuristics. These are simple, exact methods which suit relatively simple problems. Common examples include reorder point heuristics (Federgruen and Zheng, 1992; Silver et al., 1998) or the Economic Order Quantity model (Erlenkotter, 1990). However, a key limitation of these heuristics is their lack of adaptability. They rely on pre-defined parameters that may not be effective in dynamic environments with fluctuating demand or unforeseen circumstances.
Moreover, inventory management problems can also be tackled using optimization techniques like Linear Programming (LP) (Janssens and Ramaekers, 2011) or Mixed-Integer Linear Programming (MILP) (You and Grossmann, 2008). LP is well-suited for problems with linear relationships, while MILP offers more flexibility by incorporating integer variables to address specific complexities like minimum order quantities. However, these methods do not directly account for uncertainty.
If the optimal policy is analytically intractable, methods like stochastic optimization can be used to find an approximate optimal policy (Grossmann et al., 2016; You and Grossmann, 2011). While stochastic optimization often finds good solutions, it does not guarantee optimality. In specific scenarios, exact numerical methods like dynamic programming might be applicable (Berovic and Vinter, 2004; Perez et al., 2021). However, these methods are often limited by scale due to computational demands. This leads to approximate numerical methods. These methods provide scalability but may be at a sub-optimal performance. Additionally, Approximate Dynamic Programming (ADP) is a powerful technique specifically designed for problems with both dynamics (e.g., decisions impact future states) and stochasticity (e.g., uncertain outcomes) (Katanyukul et al., 2011). ADP leverages approximations and sampling techniques to find good solutions for complex inventory management problems that might be intractable for traditional methods. Reinforcement Learning methods, closely related to dynamic programming, also belong to the approximate numerical methods class.
Reinforcement learning has emerged as a promising approach to address the dynamic and complex nature of inventory management systems (Giannoccaro and Pontrandolfo, 2002; Bharti et al., 2020a; Boute et al., 2022; Rangel-Martinez and Ricardez-Sandoval, 2024). Single-Agent RL methodologies have been explored extensively in this context to solve a Markov Decision Process (MDP), the mathematical framework used to model decision making in an inventory management system. Unlike traditional heuristic approaches that rely on pre-defined rules, RL leverages interaction with the environment to learn optimal policies. Reinforcement learning, particularly deep reinforcement learning, leverages neural networks to approximate the value functions and policies, enabling it to handle high-dimensional state and action spaces. Methods like Q-learning (Chaharsooghi et al., 2008; Kara and Dogan, 2018; Bharti et al., 2020b; Vera, 2021), Deep Q-Networks (DQN) (Oroojlooyjadid et al., 2022; Meisheri et al., 2020), and Policy Gradient methods (Stranieri and Stella, 2022; Siems et al., 2023; Burtea and Tsay, 2024) have shown promise in developing adaptive and scalable inventory policies that can learn from interactions with the environment over time. However, RL often struggles with integer or mixed-integer decisions, which are common in inventory management problems (e.g., order quantities). As the scale of inventory systems increases, the complexity of the decision-making process grows, resulting in a significantly larger action space. This increased complexity can hinder the effectiveness of RL algorithms, as they require extensive training data and longer convergence times to identify optimal policies. This limitation is particularly relevant because studies using RL for inventory control often represent the act of order replenishment as a discrete integer decision Chaharsooghi et al. (2008); Stranieri and Stella (2022), rather than leveraging pre-defined heuristic parameters like reorder point (s) or order-up-to level (S) which could speed up adoption in industry.
There has been a growing interest to leverage the graph structure of a supply chain by using Graph Neural Networks (GNNs). The main idea is to leverage GNNs to learn the hidden representation of the data encoded as a graph structure (Munikoti et al., 2022). This allows RL agents to efficiently adapt to changes in the problem domain. The application of GNNs to RL has been widely studied for a series of well known combinatorial optimization problems (Khalil et al., 2017; Mazyavkina et al., 2021) such as the vehicle routing problem (Munikoti et al., 2022), travel salesman problem (Munikoti et al., 2022) and the job shop scheduling problem (Zhang et al., 2020a, b). GNNs have been shown to improve performance across different graph sizes and types. These works provide an initial foundation but they consider simplified supply chain instances with deterministic lead times and assume centralized information. Unfortunately, real-world supply chains face inherent information sharing constraints. These constraints can arise due to technical limitations caused by incompatible systems, the inherent complex nature of supply chain structures, or even due to privacy concerns about sharing sensitive data.
@@ 1.4 Related Work
To address these limitations and leverage the collaborative nature of supply chains, multi-agent RL approaches have gained significant momentum but remain limited. These approaches model multiple agents, each representing a decision-maker within the supply chain, and allow them to interact and learn from each other. The application of MARL to the inventory management problem is relatively limited compared to other domains including single-agent RL. However, there is growing recognition of the potential benefits of MARL in addressing the dynamic and collaborative nature of inventory management in multi-agent environments. Liu et al. (2022b) applied Heterogeneous-Agent Proximal Policy Optimization (HAPPO) to a serial supply chain which showed overall better performance than single-agent RL. The study also concluded that information sharing between entities helps alleviate the bullwhip effect. Another study conducted by (Feng et al., 2022b), applied a Decentralized PPO framework on a single store problem with a large number of stock keeping units. This approach accelerated policy learning compared to standard MARL algorithm. Additionally, (Khirwar et al., 2023) applied a PPO variation MARL framework to a multi-echelon environment, and it outperformed other base-stock policies. Sultana et al. (2020), applied a multi-agent advantage actor critic algorithm to a multi-echelon, multi-product system but assumed a lead time of zero. Finally, Mousa et al. (2023) provided an analysis of multi-agent reinforcement learning algorithms for decentralised inventory management systems. The study showed that MAPPO outperformed other MARL methods and further highlighted MARL as a promising decentralized control solution for large-scale stochastic systems. While these studies presented in inventory management have made progress in applying MARL techniques, there are still opportunities for improvement. One area deserving further attention is how to leverage the inherent graph structure of a supply chain to enhance RL frameworks’ performance. By exploiting the connectivity and dependencies between different supply chain nodes, we can develop more effective algorithms that address the complexities of multi-agent inventory management.
@@ 1.5 Motivation
In this work, we propose leveraging the capabilities of GNNs to learn hidden representations of agents and their interactions. These representations are then integrated into a MARL framework to find optimal inventory policies in a multi-echelon supply chain network. Our contributions are summarized as follows:
@@
    We leverage the inherent graph structure of a supply chain to aid collaboration between entities in a supply chain.
@@
    We reduce dimensionality and increase scalability of a MARL-GNN framework by incorporating a global mean pooling aggregation mechanism within our algorithmic framework.
@@
  • 5.
    We demonstrate the effectiveness of our method in terms of scalability through four multi-echelon supply chain scenarios.

### v2

@@ Abstract
Our approach redefines the action space by parameterizing heuristic inventory control policies, into an adaptive, continuous form where parameters dynamically adjust based on system conditions and avoid combinatorial explosion typical of discrete actions. By leveraging the inherent graph structure of supply chains, our framework enables agents to learn the system’s topology, and we employ a centralized learning, decentralized execution scheme that allows agents to learn collaboratively while overcoming information-sharing constraints. Additionally, we incorporate global mean pooling and regularization techniques to enhance performance.
@@ 1.1 Inventory Control
The theory of inventory control can be traced back to the news-vendor problem (Edgeworth, 1888; Clark and Scarf, 1960), and the first widely used numerical solutions for inventory optimization seems to be the Economic Order Quantity (EOQ) model (Erlenkotter, 1990). These works were fundamental to the widely known policies in inventory control theory: (R,S), (s,S), (R,s,S) and (R,Q) (Federgruen and Zheng, 1992; Silver et al., 1998). Over the years, control theory has had an influence on production planning and inventory control (Aggarwal, 1974; Jackson et al., 2020).
One approach to find optimal inventory control policies is using traditional heuristics. These are simple, exact methods which suit relatively simple problems. Methods such as reorder point heuristics (Federgruen and Zheng, 1992; Silver et al., 1998) and the Economic Order Quantity model (Erlenkotter, 1990), provide exact solutions for structured settings. However, they lack adaptability, as they depend on pre-defined parameters that may not be effective in dynamic environments with fluctuating demand or unforeseen circumstances.
The complexity of modern systems has driven the adoption of dynamic programming, which enables the formulations of optimal policies by considering future states. However, obtaining exact analytical solutions often proves infeasible due to computational demands, especially in large-scale or highly stochastic environments.(Bellman, 1952). This limitation has paved the way for advanced numerical and optimization-based techniques beyond traditional closed-form solutions. For example, optimization techniques like Linear Programming (LP) (Janssens and Ramaekers, 2011) or Mixed-Integer Linear Programming (MILP) (You and Grossmann, 2008) are applied to inventory management problems. LP is well-suited for problems with linear relationships, while MILP offers more flexibility by incorporating integer variables to address specific complexities like minimum order quantities. However, these methods do not directly account for uncertainty, making them less effective in highly stochastic environments.
To improve adaptability in dynamic settings, feedback-based control strategies such as Model Predictive Control (MPC) have been employed. MPC uses a rolling-horizon optimization framework, incorporating real-time system updates to enhance decision-making. Unlike static optimization, MPC continuously updates decisions based on new information, making it particularly useful for inventory systems with demand variability (Dong et al., 2012). One of the key advantages of MPC is its capacity to optimize control actions by predicting future system behavior over a finite horizon. This predictive capability enables the formulation of control strategies that can respond to changes in system dynamics, thereby improving decision-making processes in inventory management (Liu et al., 2011). For instance, Liu et al. (2011) demonstrated the application of MPC in optimizing production and distribution systems, highlighting its effectiveness in handling dynamic inventory challenges. Furthermore, the integration of feedback mechanisms allows MPC to adjust its predictions based on actual system performance, thereby reducing the impact of demand variability (Conte and Pennesi, 2005). However, its computational complexity scales poorly and it does not explicitly account for uncertainty, limiting its feasibility in large, stochastic environments (Ghaemi et al., 2007).
In contrast, stochastic (Küçükyavuz, 2011) and distributionally-robust optimization approaches (Bertsimas et al., 2019; Qiu et al., 2021), offer ways to handle uncertainty explicitly. However, they face challenges such as online tractability, limiting their practical applicability in real-world supply chain scenarios. If the optimal policy is analytically intractable, methods like stochastic optimization can be used to find an approximate optimal policy (Grossmann et al., 2016; You and Grossmann, 2011). While stochastic optimization often finds good solutions, it does not guarantee optimality. In specific scenarios, exact numerical methods like dynamic programming might be applicable (Berovic and Vinter, 2004; Perez et al., 2021). However, these methods are often limited by scale due to computational demands. This leads to approximate numerical methods. These methods provide scalability but may be at a sub-optimal performance. Additionally, Approximate Dynamic Programming (ADP) is a powerful technique specifically designed for problems with both dynamics (e.g., decisions impact future states) and stochasticity (e.g., uncertain outcomes) (Katanyukul et al., 2011). ADP leverages approximations and sampling techniques to find good solutions for complex inventory management problems that might be intractable for traditional methods. Reinforcement Learning methods, closely related to dynamic programming, also belong to the approximate dynamic programming methods class.
@@ 1.2 Reinforcement learning for Inventory Control
Reinforcement learning (RL) has emerged as a promising alternative for addressing the challenges of stochastic sequential decision-making problems. Its ability to excel in complex, dynamic environments and handle uncertainty, places RL as a valuable tool to enhance decision making in supply chains. While closely linked to dynamic programming, RL offers a general solution to identify approximate optimal policies for stochastic processes by leveraging the Bellman optimality equation to iteratively update value functions and improve decision-making policies over time. RL also provides a cost-effective solution for decision-making by enabling offline training, which reduces the online computational overhead compared to optimization approaches that require continual updates in receding or shrinking horizon frameworks.
Single-Agent RL methodologies have been explored extensively in this context to solve a Markov Decision Process (MDP), the mathematical framework used to model decision making in an inventory management system. Unlike traditional heuristic approaches that rely on pre-defined rules, RL leverages interaction with the environment to learn optimal policies. Reinforcement learning, particularly deep reinforcement learning, leverages neural networks to approximate the value functions and policies, enabling it to handle high-dimensional state and action spaces. Methods like Q learning (Chaharsooghi et al., 2008; Kara and Dogan, 2018; Bharti et al., 2020; Vera, 2021) and Policy Gradient methods (Siems et al., 2023; Burtea and Tsay, 2024; Rangel-Martinez and Ricardez-Sandoval, 2023, 2024; Shin et al., 2019; Yoo et al., 2021)have shown promise in developing adaptive and scalable inventory policies that can learn from interactions with the environment over time. However, RL often struggles with integer or mixed-integer decisions, which are common in inventory management problems (e.g., order quantities). As the scale of inventory systems increases, the complexity of the decision-making process grows, resulting in a significantly larger action space. This increased complexity can hinder the effectiveness of RL algorithms, as they require extensive training data and longer convergence times to identify optimal policies.
To address these challenges, recent work has explored action parametrization methods that enable RL to manage complex or hierarchical action spaces by combining continuous and discrete decision variables (Fan et al., 2019; Bester et al., 2019). In the context of multi-agent control and decision-making, Fan et al. (2019) introduced a hybrid actor-critic reinforcement learning model that operates in a parameterized action space, allowing the agent to handle both discrete and continuous action variables simultaneously. Their approach was demonstrated in RL environments such as Catching Point, Moving, Chase and Attack, and Half Field Football. Khamassi and Tzafestas (2016) and Zhang et al. (2024) introduced Parameterized Action Space Markov Decision Processes (PAMDPs) in the context of robotics and other RL environments, which demonstrated how continuous action spaces can be effectively utilized in complex manipulation tasks. This is relevant in inventory management as decisions often involve selecting order quantities which can be treated as discrete actions with continuous parameters. However, current studies in inventory control often represent the act of order replenishment as a discrete integer decision Chaharsooghi et al. (2008); Stranieri and Stella (2022), rather than leveraging on pre-defined heuristic parameters like reorder point (s) or order-up-to level (S).. Notably, the action parametrization of dynamic parameters for such heuristic policies, (e.g., dynamical tuning $(s,S)$ values) has not yet ben extensively explored. This gap in the literature opens the door for RL methods that dynamically optimize these parameters, potentially enabling faster adoption in industry by blending RL’s adaptability with the structure of traditional heuristics.
However, RL’s effectiveness relies on the availability of information sharing among the entities within the supply chain where individual actors must collaborate under uncertainty and coordination constraints. In scenarios where multiple interconnected entities are involved, ensuring seamless information exchange can be challenging due to factors such as data privacy concerns, proprietary information, or communication constraints. This limitation may hinder the adoption and implementation of RL-based approaches in complex supply chain environments, highlighting the need for strategies to address information sharing challenges effectively.
Figure 1: Multi-Agent Reinforcement Learning with Graph Neural Networks for Inventory Management – A decentralized policy learning approach where each warehouse optimizes its inventory decisions using independent neural network policies $\pi_{1},\pi_{2},\pi_{3}$ while a Graph Neural Network (GNN) captures spatial dependencies between them, enabling coordinated decision-making across the supply chain.
@@ 1.3 Multi-Agent Reinforcement Learning
Multi-Agent Reinforcement Learning (MARL) provides a framework for addressing such challenges by enabling decentralized decision-making across multiple entities. MARL frameworks allow multiple entities within a supply chain to autonomously learn and adapt their decision-making processes while interacting with each other. By leveraging MARL, supply chain entities can collaborate and coordinate their actions more effectively, leading to improved overall performance and resilience against disruptions. Additionally, MARL provides a scalable approach to handle the complexities of large-scale supply chain systems by distributing decision-making across multiple agents.
@@
The application of MARL to the inventory management problem is relatively limited compared to other domains including single-agent RL. However, there is growing recognition of the potential benefits of MARL in addressing the dynamic and collaborative nature of inventory management in multi-agent environments. Liu et al. (2022c) applied Heterogeneous-Agent Proximal Policy Optimization (HAPPO) to a serial supply chain which showed overall better performance than single-agent RL. The study also concluded that information sharing between entities helps alleviate the bullwhip effect. Another study conducted by (Feng et al., 2022b), applied a Decentralized PPO framework on a single store problem with a large number of stock keeping units. This approach accelerated policy learning compared to standard MARL algorithm. Additionally, (Khirwar et al., 2023) applied a PPO variation MARL framework to a multi-echelon environment, and it outperformed other base-stock policies. Sultana et al. (2020), applied a multi-agent advantage actor critic algorithm to a multi-echelon, multi-product system but assumed a lead time of zero. Finally, Mousa et al. (2023) provided an analysis of multi-agent reinforcement learning algorithms for decentralised inventory management systems. The study showed that MAPPO outperformed other MARL methods and further highlighted MARL as a promising decentralized control solution for large-scale stochastic systems.
However, despite its potential, MARL faces several challenges. One issue is the non-stationarity of the learning environment inherent in MARL (Tan, 1993; Tampuu et al., 2017). This occurs when multiple agents are learning simultaneously so the transition dynamics are not stationary. One setting commonly used to overcome this is to use a central critic during training that has access to global observations and actions (Nekoei et al., 2023). This is known as the Centralized Training, Decentralized Execution learning paradigm. However, while adding more information can help mitigate the non-stationarity problem, it can also lead to new issues. Naively concatenating all available information can result in information overload and inefficiencies (Lowe et al., 2017; Yu et al., 2022; Nayak et al., 2023). This indiscriminate accumulation of data can cause policy overfitting, where agents develop strategies that perform well on the excessive training information but poorly in real-world scenarios due to the lack of generalization (Nayak et al., 2023; Hu et al., 2021). Therefore, the development of novel techniques for smart information aggregation is required to avoid policy overfitting and enhance the efficacy of MARL frameworks in supply chain management.
Traditional MARL methods also lack structural awareness as they treat agent interactions as homogeneous and symmetric, failing to account for the hierarchical and graph-based structure often inherent in systems like supply chains. By failing to capture such dependencies, these methods struggle to model the nuanced coordination required in multi-entity systems. Coordination challenges are further highlighted under partial information settings, when privacy concerns or communication constraints limit the availability of data. Many MARL approaches assume full observability or reliable communication, making them ill-suited for real-world supply chains, where information exchange is often fragmented or restricted. This limitation highlights the need for methods that can operate effectively under incomplete information while leveraging structural insights to enhance coordination.
These challenges highlight the need for approaches that can leverage the inherent structure and connectivity of supply chains. Graph-based methods offer a compelling solution by explicitly modeling the hierarchical relationships and dependencies between agents. Representing supply chains as graph-structured systems—where nodes correspond to entities such as suppliers, warehouses, and retailers, and edges denote interactions like material flows or shared information—allows MARL frameworks to better capture the dynamics of multi-agent coordination.
There has been a growing interest to leverage the graph structure of a supply chain by using Graph Neural Networks (GNNs). The main idea is to leverage GNNs to learn the hidden representation of the data encoded as a graph structure (Munikoti et al., 2022). This allows RL agents to efficiently adapt to changes in the problem domain. The application of GNNs to RL has been widely studied for a series of well known combinatorial optimization problems (Khalil et al., 2017; Mazyavkina et al., 2021) such as the vehicle routing problem (Munikoti et al., 2022), travel salesman problem (Munikoti et al., 2022)and the job shop scheduling problem (Zhang et al., 2020a, b). GNNs have been shown to improve performance across different graph sizes and types. These works provide an initial foundation but they consider simplified supply chain instances with deterministic lead times and assume centralized information. However, real-world supply chains face inherent information sharing constraints that can arise due to technical limitations caused by incompatible systems, the inherent complex nature of supply chain structures, or even due to privacy concerns about sharing sensitive data. Therefore, there is a need to develop decentralized decision-making frameworks that not only overcome information-sharing constraints but also effectively incorporate the hierarchical and graph-based structures inherent in supply chains
@@ 1.4 Motivation
In this work, we propose leveraging the capabilities of GNNs to learn spatial representations of agents and their interactions. These representations are then integrated into a MARL framework to find optimal inventory policies in a multi-echelon supply chain network. Our contributions are summarized as follows:
@@
    We leverage the inherent graph structure and geometric properties of a supply chain to aid collaboration between entities in a supply chain.
@@
    We reduce dimensionality and increase scalability of the MARL-GNN framework by incorporating a global mean pooling aggregation mechanism within our algorithmic framework.
@@ 2 Preliminaries
2.1 Markov Decision Process (MDP)
A Markov Decision Process (MDP) is a mathematical framework used to describe a sequential decision-making problem for a single agent interacting with an environment. An MDP is defined as a tuple $\left\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\right\rangle$ where $\mathcal{S}$ is the set of all states, $\mathcal{A}$ is the set of actions, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the state transition probability function where $\mathcal{T}(s^{\prime}|s,a)$ defines the probability of transitioning to a state $s^{\prime}\in\mathcal{S}\subseteq\mathbb{R}^{n_{s}}$ given that the agent is currently in state $s\in\mathcal{S}\subseteq\mathbb{R}^{n_{s}}$ and takes action $a\in\mathcal{A}\subseteq\mathbb{R}^{n_{a^{\prime}}}$. For a deterministic policy $\pi$, the agent takes actions $a_{t}=\pi(s_{t})$, while for a stochastic policy, the action is sampled from a policy $\pi$ represented by a conditional probability distribution $a_{t}\sim\pi(\cdot|s_{t})$. $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function which specifies the immediate reward $r_{t}$ received after transitioning from state $s$ to $s^{\prime}$ by taking action $a$ and $\gamma\in[0,1]$ is the discount factor. For further treatment of the subject, readers are referred to Sutton (2018).